# DynoSurf: Neural Deformation-based Temporally Consistent Dynamic Surface Reconstruction

### v1

@@ 3.3 Temporal Reconstruction via Learnable Deformation Field
Control points blending-based learnable deformation field.  We first build a set of control point $\mathbf{U}:=\{\mathbf{u}_{1},...,\mathbf{u}_{|\mathbf{U}|}\}$, which are obtained using the farthest point sampling (FPS) algorithm eldar1997farthest on the vertex set of the tetrahedron mesh $\mathbb{T}$ (i.e., $\mathbf{Q}$) such that the control points $\mathbf{u}_{r}$, $r=1,2,\dots,|\mathbf{U}|$ are uniformly distributed in $\mathbb{T}$. Note that the positions of initial control points are also adaptively optimized during training to make the deformation more flexible. We associate each $\mathbf{u}_{r}$ with a rotation matrix $\mathbf{R}_{r}^{k}\in\texttt{SO}(3)$ and a translation $\mathbf{t}_{r}^{k}\in\mathbb{R}^{3}$ for each frame $\mathbf{P}_{k}$, $k=1,2,\dots,K$. Then for each vertex of the template surface $\mathbf{\overline{v}}\in\overline{\mathbf{V}}$, which is the vertex set of the surface $\overline{\mathbb{M}}$. the deformed vertex at the $k$-th frame is represented as
@@
where $\mathbf{F}_{\bf{\Omega}}(\cdot)$ is an MLP parameterized with $\bf{\Omega}$, and [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the concatenation operator. We initialize $\omega(\mathbf{\overline{v}},\mathbf{u}_{r})$ as
@@
where $\eta=0.1$ by default. The reconstructed surface of $\mathbf{P}_{k}$ can be obtained as $\mathbb{M}_{k}=\{\mathbf{V}_{k},\mathcal{E}\}$ with $\mathbf{V}_{k}$ being the collection of $\mathbf{v}^{k}$.
We also utilize an MLP $F_{\bf{\Xi}}(\cdot)$ parametrized with $\bf{\Xi}$ to learn the deformation transformations of all frames, i.e.,
$[\xi_{r}^{k},\mathbf{t}_{r}^{k}]=F_{\bf{\Xi}}(\Gamma([\mathbf{u}_{r}||k])),$ (9)
where $\xi_{r}^{k}\in\texttt{so}(3)$ is the Lie algebra format of $\mathbf{R}_{r}^{k}$. Owing to the spectral bias property of neural networks rahaman2019spectral , the learned transformations would vary smoothly in both spatial and temporal domains to ensure the spatial and temporal smoothness of the constructed mesh sequence.
Remark. It’s worth noting that our deformation formula is different from the neural deformation graph in bozic2021neural , which only defines the importance weight for each graph node, and blends them according to the spatial positions. Here, we use an MLP to parameterize the weights of any point in space. Additionally, it confines the node positions within the object. These constraints limit the flexibility of the deformation. The surface deformation in RMA-Net feng2021recurrent is defined based on globally shared transformations, which limits the ability to capture the locally correlated characteristics of non-rigid object deformations. By using control points-blending, we can effectively enhance the representation capability of the deformation.
@@
$\footnotesize{L_{\textrm{R-CD}}^{k}=\frac{1}{|\widetilde{\mathbf{V}}_{k}|}\sum_{\mathbf{v}_{i}\in\widetilde{\mathbf{V}}_{k}}\phi_{\alpha}({\mathbf{v}}_{i},\mathbf{p}_{\rho_{i}})\cdot\|{\mathbf{v}}_{i}-\mathbf{p}_{\rho_{i}}\|^{2}+\frac{1}{|\mathbf{P}_{k}|}\sum_{\mathbf{p}_{j}\in\mathbf{P}_{k}}\phi_{\alpha}(\mathbf{p}_{j},{\mathbf{v}}_{\tau_{j}})\cdot\|\mathbf{p}_{j}-{\mathbf{v}}_{\tau_{j}}\|^{2},}$ (11)
where $\widetilde{\mathbf{V}}_{k}$ is the set of points uniformly sampled from $\mathbb{M}_{k}$ with the size of $10^{4}$ by default, and $\phi_{\alpha}(x,y)=\exp(-\alpha\cdot\|x-y\|^{2})$. The normal loss term $L_{\textrm{Norm}}=\sum_{k}\texttt{NC}_{\ell_{1}}(\widetilde{\mathbf{V}}_{k},\mathbf{P}_{k})/K$ constrains the consistency of normals.
The proposed robust SDF loss term $L_{{\textrm{R-SDF}}}$ regularizes the orientation of the deformed mesh, defined as
@@
where $\widehat{\mathbf{q}}^{k}$ is the deformed positions of $\mathbf{q}$ according to Eq. (6); $\overline{\psi}_{\beta}^{k}(\mathbf{q})=\frac{\psi_{\beta}^{k}(\mathbf{q})}{\sum_{k}\psi_{\beta}^{k}(\mathbf{q})}$ with ${\psi}_{\beta}^{k}(\mathbf{q})=\exp(-\beta\cdot|\texttt{SDF}_{{\rm IMLS}}(\widehat{\mathbf{q}}^{k},\mathbf{P}_{k})|^{2})$ and $\beta=50$ is to measure the confidence of each term. That is, when the deformed point $\widehat{\mathbf{q}}^{k}$ has a smaller value in the SDF approximated by the target point cloud, it is closer to the surface, and we assume that it is more accurate; $\mathbf{Q}_{s}$ is a subset of size $|\mathbf{Q}_{s}|=10^{4}$ randomly sampled from $\mathbf{Q}$. $f_{\gamma}(x)=\frac{1}{1+\exp({-\gamma\cdot x})}$ with $\gamma=10^{2}$ is used to scale the amplified SDF value to $(0,1)$ because we pay more attention to whether the point $\widehat{\mathbf{q}}^{k}$ is inside or outside the surface.
@@
where $\mathcal{I}(k)=\{k-1,k+1\}$. When $k=1$ or $k=K$, we only use one neighbor point cloud frame. Finally, it is crucial to ensure that the enhanced template surface maintains a reasonable shape throughout the joint optimization process. Thus, we constrain the deformation field of the template surface as unit transformation as possible with a loss term $L_{\textrm{Shape}}$:
@@ 4.1 Experimental Settings
Implementation details.  In our approach, we employed Multi-Layer Perceptrons (MLPs) consisting of 5 linear layers with a feature dimension of 128 to implement $F_{\bf{\Phi}}(\cdot)$, $F_{\bf{\Xi}}(\cdot)$, and $F_{\bf{\Omega}}(\cdot)$. The default weights setting in the loss function are listed in Tab. 1. Training was conducted using the ADAM optimizer kingma2014adam with a learning rate of $1\times 10^{-4}$. We performed all experiments on a single NVIDIA A6000 GPU.
Table 1: Default parameter settings and runtime. For the learning template surface, we listed iterations and running time for coarse/fine stage.
Learning template surface Temporal reconstruction
500 0.001 50 $10^{3}$/5000 10 s/2 mins 500 0.001 100 $10^{3}$ 1 $10^{4}$ 30 mins
Datasets.  We utilized three well-established benchmark datasets, namely DFA-UST bogo2017dynamic , DT4D li20214dcomplete , and AMA vlasic2008articulated , for our evaluation. From motion sequences, we selected 17 consecutive frames to form a set of input point clouds. Randomly sampling 5000 points from the ground truth mesh, we ensured diversity in our point cloud selection. Adopting the train and test data splits provided by lei2022cadex , we used 109 sets and 89 sets from the test data of the DFAUST dataset and DT4D-animal dataset, respectively. The AMA dataset comprises 10 motion sequences; for supervised learning comparison methods, we employed 1176 sets from 7 sequences as the train set and 32 sets of point clouds from the remaining 3 sequences ("crane", "march1" and "samba") as the test set. To maintain consistency across sequences, we normalized the diagonal length of the bounding box in the first frame to 1 and centered it at the origin of the coordinates. Subsequent point clouds underwent transformations based on the scale and translation of the first frame. We utilized Meshlab’s point cloud estimation normal method for normal estimation and adjusted the normal directions to face outward.
Refer to caption(a) Input(b) LPDC tang2021learning (c) Cadex lei2022cadex (d) Ours00.06
@@
DFAUST bogo2017dynamic LPDC tang2021learning 2.430 0.929 0.299 0.633 1.46
Cadex lei2022cadex 1.062 0.941 0.519 0.823 1.28
@@
AMA vlasic2008articulated LPDC tang2021learning 108.822 0.666 0.045 0.101 14.1
Cadex lei2022cadex 71.041 0.663 0.052 0.119 13.6
Ours 0.320 0.918 0.943 0.683 4.44
Refer to caption(a) Input(b) LPDC tang2021learning (c) Cadex lei2022cadex (d) Ours00.15
Figure 6: Comparison of visual results by different methods on the AMA dataset
Methods under comparison.  We conducted a comparative analysis involving our proposed method and two cutting-edge techniques specialized in dynamic surface reconstruction from sequences of point clouds: LPDC tang2021learning and Cadex lei2022cadex . Note that both Cadex and LPDC necessitate ground-truth occupancy values and ground-truth correspondences for supervision during training. For Cadex, we employed the officially released pre-trained model provided by the authors for testing on DFAUST and DT4D datasets. As for LPDC, the authors exclusively performed experiments on the DFAUST dataset and released the pre-trained model. To ensure a comprehensive comparison, we introduced the AMA dataset, a human body dataset characterized by significantly larger deformation and movement ranges compared to DFAUST. Furthermore, we retrained Cadex and LPDC on the AMA dataset.
Evaluation metrics.  Following previous works mescheder2019occupancy ; lei2022cadex ; peng2021shape , we scaled the test data to its original size and quantitatively assessed various methods using $\ell_{2}$-norm-based Chamfer Distance (CD), Normal Consistency (NC), F-score with thresholds of $0.5\%$ (F-$0.5\%$) and $1\%$ (F-$1\%$), and Correspondences Error (Corr.) to evaluate different methods quantitatively. For the visual results, we assigned the same colors to corresponded vertices to indicate the accuracy of learned correspondences by different methods. In addition, we also visualized the pointwise distance error from the ground-truth surface.
@@ 4.2 Comparisons with State-of-the-Art Methods
Results of noisy and partially missing data.  We also evaluate the performance of our method on noisy data. We add $0.5\%$ Gaussian noise for each point in a point cloud sequence in DFAUST dataset bogo2017dynamic and show the results in Fig. 7. Since point cloud data collected by real-depth cameras often contain missing parts, we tested the situation when there are some holes in the surface. For a sequence in "swing" motion sequences of AMA dataset vlasic2008articulated , we placed two virtual cameras in front of and behind the mesh model for simulated acquisition and merged the point clouds scanned by the two cameras. Fig. 7 shows the performance of our method and the variants. We can see that our method can reconstruct partially missing meshes. More results are presented in the Supplementary Material.
@@ 4.3 Ablation Study
Table 3: Ablation studies of our method on the AMA dataset vlasic2008articulated . The best and second best results are highlighted in bold and underlined.
@@
MLP-Deform 0.412 0.909 0.917 0.618 5.91
w. $\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})$ 0.475 0.911 0.890 0.564 3.26
Fixed $\{\mathbf{u}_{r}\}$ 0.372 0.911 0.927 0.637 3.23
w.o. $L_{\textrm{R-SDF}}$ 0.330 0.918 0.937 0.649 3.14
w.o. $L_{\textrm{Norm}}$ 0.656 0.900 0.884 0.582 3.30
w.o. $L_{\textrm{Smo}}$ 0.378 0.914 0.923 0.616 3.30
w.o. $L_{\textrm{Shape}}$ 0.335 0.919 0.936 0.636 2.91
$|\mathbf{U}|=10$ 0.437 0.910 0.910 0.594 3.30
$|\mathbf{U}|=30$ (Ours) 0.324 0.919 0.939 0.649 2.89
$|\mathbf{U}|=100$ 0.315 0.919 0.943 0.651 3.16
@@
  • Fixed blending weight formula: We replaced the learnable blending weights in Eq. (6) by setting it to the value computed by $\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})$ in Eq. (8) throughout the learning process.
@@
3) Loss functions during temporal reconstruction.  To assess the effectiveness of various loss terms used in the temporal reconstruction stage, we individually excluded the robust SDF loss $L_{\textrm{R-SDF}}$, deformation smooth loss $L_{\textrm{Smo}}$, and shape maintenance loss $L_{\textrm{Shape}}$, while keeping other aspects unchanged. The results presented in Tab. 3 demonstrate their necessity and contributions.
4) Keyframe selection and template surface learning.  Compared to directly setting the middle frame as a keyframe, Fig. 10 shows that the selected keyframe by our strategy is closer to the average shape. After learning the template surface, we also show the results of only using the fine stage (Fig. 10 (b)) and using the coarse-to-fine stage (Fig. 10 (c)). Based on Fig. 10 (c), we further show the enhanced template surface after the temporal reconstruction process.

### v2

@@ 3.3 Temporal Reconstruction via Learnable Deformation Field
Control points blending-based learnable deformation field.  We first build a set of control point $\mathbf{U}:=\{\mathbf{u}_{1},...,\mathbf{u}_{|\mathbf{U}|}\}$, which are obtained using the farthest point sampling (FPS) algorithm [11] on the vertex set of the tetrahedron mesh $\mathcal{T}$ (i.e., $\mathbf{Q}$) such that the control points $\mathbf{u}_{r}$, $r=1,2,\dots,|\mathbf{U}|$ are uniformly distributed in $\mathcal{T}$. Note that the positions of initial control points are also adaptively optimized during training to make the deformation more flexible. We associate each $\mathbf{u}_{r}$ with a rotation matrix $\mathbf{R}_{r}^{k}\in\texttt{SO}(3)$ and a translation $\mathbf{t}_{r}^{k}\in\mathbb{R}^{3}$ for each frame $\mathbf{P}_{k}$, $k=1,2,\dots,K$. Then for each vertex of the template surface $\mathbf{\overline{v}}\in\overline{\mathbf{V}}$, which is the vertex set of the surface $\overline{\mathcal{M}}$. The deformed vertex at the $k$-th frame is represented as
@@
where ${F}_{\bf{\Omega}}(\cdot)$ is an MLP parameterized with $\bf{\Omega}$, and [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the concatenation operator. Before training, we pre-optimize $\bf{\Omega}$ so that $\omega(\mathbf{\overline{v}},\mathbf{u}_{r})=\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})$, where
@@
The reconstructed surface of $\mathbf{P}_{k}$ can be obtained as $\mathcal{M}_{k}=\{\mathbf{V}_{k},\mathbf{F}\}$ with $\mathbf{V}_{k}$ being the collection of $\mathbf{v}^{k}$.
We also utilize an MLP $F_{\bf{\Xi}}(\cdot)$ parametrized with $\bf{\Xi}$ to learn the deformation transformations of all frames, i.e., $[\xi_{r}^{k},\mathbf{t}_{r}^{k}]=F_{\bf{\Xi}}(\Gamma([\mathbf{u}_{r}||k]))$, where $\xi_{r}^{k}\in\texttt{so}(3)$ is the Lie algebra format of $\mathbf{R}_{r}^{k}$. Owing to the spectral bias property of neural networks [43], the learned transformations would vary smoothly in both spatial and temporal domains to ensure the spatial and temporal smoothness of the constructed mesh sequence.
@@
$L_{\textrm{R-CD}}^{k}=\frac{1}{|\widetilde{\mathbf{V}}_{k}|}\sum_{\mathbf{v}_{i}\in\widetilde{\mathbf{V}}_{k}}\phi_{\alpha}({\mathbf{v}}_{i},\mathbf{p}_{\rho_{i}})\cdot\|{\mathbf{v}}_{i}-\mathbf{p}_{\rho_{i}}\|^{2}+\frac{1}{|\mathbf{P}_{k}|}\sum_{\mathbf{p}_{j}\in\mathbf{P}_{k}}\phi_{\alpha}(\mathbf{p}_{j},{\mathbf{v}}_{\tau_{j}})\cdot\|\mathbf{p}_{j}-{\mathbf{v}}_{\tau_{j}}\|^{2},$ (9)
where $\widetilde{\mathbf{V}}_{k}$ is the set of points uniformly sampled from $\mathcal{M}_{k}$, and $\phi_{\alpha}(x,y)=\exp(-\alpha\cdot\|x-y\|^{2})$. $\mathbf{p}_{\rho_{i}}$ and $\mathbf{v}_{\tau_{j}}$ are the closest points for $\mathbf{v}_{i}$ and $\mathbf{p}_{j}$ respectively. The normal loss term $L_{\textrm{Norm}}=\sum_{k}\texttt{NC}_{\ell_{1}}(\widetilde{\mathbf{V}}_{k},\mathbf{P}_{k})/K$ constrains the consistency of normals. The proposed robust SDF loss term $L_{{\textrm{R-SDF}}}$ regularizes the orientation of the deformed mesh, defined as
@@
where $\widehat{\mathbf{q}}^{k}$ is the deformed position of $\mathbf{q}$ according to Eq. (5); $\overline{\psi}_{\beta}^{k}(\mathbf{q})=\frac{\psi_{\beta}^{k}(\mathbf{q})}{\sum_{k}\psi_{\beta}^{k}(\mathbf{q})}$ with ${\psi}_{\beta}^{k}(\mathbf{q})=\exp(-\beta\cdot|\texttt{SDF}_{{\rm IMLS}}(\widehat{\mathbf{q}}^{k},\mathbf{P}_{k})|^{2})$ is to measure the confidence of each term. That is, when the deformed point $\widehat{\mathbf{q}}^{k}$ has a smaller value in the SDF approximated by the target point cloud, it is closer to the surface, and we assume that it is more accurate; $\mathbf{Q}_{s}$ is a subset randomly sampled from $\mathbf{Q}$. $f_{\gamma}(x)=\frac{1}{1+\exp({-\gamma\cdot x})}$ is used to scale the amplified SDF value to $(0,1)$ because we pay more attention to whether the point $\widehat{\mathbf{q}}^{k}$ is inside or outside the surface.
@@
where $\mathcal{E}$ is the edge set derived by face set $\mathbf{F}$. $\mathcal{I}(k)=\{k-1,k+1\}$. When $k=1$ or $k=K$, we only use one neighbor point cloud frame. Finally, it is crucial to ensure that the enhanced template surface maintains a reasonable shape throughout the joint optimization process. Thus, we constrain the deformation field at keyframe as close to the identity transformation as possible with a loss term $L_{\textrm{Shape}}$:
@@ 4.1 Experimental Settings
Implementation details.  In our approach, we employed MLPs consisting of 5 linear layers with a feature dimension of 128 to implement $F_{\bf{\Phi}}(\cdot)$, $F_{\bf{\Xi}}(\cdot)$, and $F_{\bf{\Omega}}(\cdot)$. Training was conducted using the ADAM optimizer [20] with a learning rate of $1\times 10^{-4}$. The Supplementary Material provides more details. We performed all experiments on a single NVIDIA A6000 GPU.
Datasets.  We utilized three well-established benchmark datasets, namely DFA-UST [4], DT4D [27], and AMA [52], for our evaluation. From motion sequences, we selected 17 consecutive frames to form a set of input point clouds. Randomly sampling 5000 points from the ground truth mesh, we ensured diversity in our point cloud selection. Adopting the train and test data splits provided by [23], we used 109 sets and 89 sets from the test data of the DFAUST dataset and DT4D-animal dataset, respectively. The AMA dataset comprises 10 motion sequences; for supervised learning comparison methods, we employed 1176 sets from 7 sequences as the train set and 32 sets of point clouds from the remaining 3 sequences ("crane", "march1" and "samba") as the test set. To maintain consistency across sequences, we normalized the diagonal length of the bounding box in the first frame to 1 and centered it at the origin of the coordinates. Subsequent point clouds underwent transformations based on the scale and translation of the first frame. We utilized the algorithm in pymeshlab [33] to estimate the normals of the point cloud and adjusted the normal directions to face outward.
Refer to caption(a) Input(b) LPDC [50](c) Cadex [23](d) Ours00.06
@@
DFAUST [4] LPDC [50] 2.430 0.929 0.299 0.633 1.46
Cadex [23] 1.062 0.941 0.519 0.823 1.28
@@
AMA [52] LPDC [50] 108.822 0.666 0.045 0.101 14.1
Cadex [23] 71.041 0.663 0.052 0.119 13.6
Ours 0.320 0.918 0.683 0.943 4.44
Refer to caption(a) Input(b) LPDC [50](c) Cadex [23](d) Ours00.15
Figure 6: Comparison of visual results by different methods on the AMA dataset.
Methods under comparison.  We conducted a comparative analysis involving our proposed method and two cutting-edge techniques specialized in dynamic surface reconstruction from sequences of point clouds: LPDC [50] and Cadex [23]. Note that both Cadex and LPDC necessitate ground-truth occupancy values and ground-truth correspondences for supervision during training. For Cadex, we employed the officially released pre-trained model provided by the authors for testing on DFAUST and DT4D datasets. As for LPDC, the authors exclusively performed experiments on the DFAUST dataset and released the pre-trained model. To ensure a comprehensive comparison, we added the comparison on the AMA dataset, a human body dataset characterized by significantly larger deformation and movement ranges compared to DFAUST. Furthermore, we retrained Cadex and LPDC on the AMA dataset.
Evaluation metrics.  Following previous works [32, 23, 40], we scaled the test data to its original size and quantitatively assessed various methods using $\ell_{2}$-norm-based Chamfer Distance (CD), Normal Consistency (NC), F-score with thresholds of $0.5\%$ (F-$0.5\%$) and $1\%$ (F-$1\%$), and Correspondences Error (Corr.) to evaluate different methods quantitatively. For the visual results, we assigned the same colors to corresponding vertices to indicate the accuracy of learned correspondences by different methods. In addition, we also show the error map to visualize the pointwise distance error from the ground-truth surface.
@@ 4.2 Comparisons with State-of-the-Art Methods
Results of noisy and partially missing data.  We also evaluate the performance of our method on noisy data. We add $0.5\%$ Gaussian noise for each point in a point cloud sequence in DFAUST dataset [4] and show the results in Fig. 7. Since point cloud data collected by real-depth cameras often contain missing parts, we tested the situation when there are some holes in the surface. For a sequence in "swing" motion sequences of AMA dataset [52], we placed two virtual cameras in front of and behind the mesh model for simulated acquisition and merged the point clouds scanned by the two cameras. Fig. 7 shows the performance of our method and the variants. We can see that our method can reconstruct partially missing meshes. More results are presented in the Supplementary Material.
@@ 4.3 Ablation Study
Table 2: Ablation studies of our method on the AMA dataset [52]. The best and second best results are highlighted in bold and underlined.
@@
MLP-Deform 0.412 0.909 0.618 0.917 5.91
w. $\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})$ 0.475 0.911 0.564 0.890 3.26
Fixed $\{\mathbf{u}_{r}\}$ 0.372 0.911 0.637 0.927 3.23
w.o. $L_{\textrm{R-SDF}}$ 0.330 0.918 0.649 0.937 3.14
w.o. $L_{\textrm{Norm}}$ 0.656 0.900 0.582 0.884 3.30
w.o. $L_{\textrm{Smo}}$ 0.378 0.914 0.616 0.923 3.30
w.o. $L_{\textrm{Shape}}$ 0.335 0.919 0.636 0.936 2.91
$|\mathbf{U}|=10$ 0.437 0.910 0.594 0.910 3.30
$|\mathbf{U}|=30$ (Ours) 0.324 0.919 0.649 0.939 2.89
$|\mathbf{U}|=100$ 0.315 0.919 0.651 0.943 3.16
@@
  • Fixed blending weight formula: We replaced the learnable blending weights in Eq. (5) by setting it to $\tilde{\omega}(\mathbf{\overline{v}},\mathbf{u}_{r})$ in Eq. (7) throughout the learning process.
@@
3) Loss functions during temporal reconstruction.  To assess the effectiveness of various loss terms used in the temporal reconstruction stage, we individually excluded the robust SDF loss $L_{\textrm{R-SDF}}$, deformation smooth loss $L_{\textrm{Smo}}$, and shape preservation loss $L_{\textrm{Shape}}$, while keeping other aspects unchanged. The results presented in Tab. 2 demonstrate their necessity and contributions.
4) Keyframe selection, template surface learning and arbitrary length of input sequences.  The performances and discussions can be found in the Supplementary Material.
5 Conclusion and Discussion
Although our method has achieved good performance on many examples, it may encounter limitations in cases where the point cloud in certain areas is sparse or the local structure is elongated but exhibits significant non-rigid deformations (as depicted on the right). This is due to our utilization of Chamfer distance as the alignment supervision, which restricts the establishment of correspondences between adjacent frames solely based on spatial proximity. Incorporating more robust metrics for shape alignment would enhance the effectiveness of our approach.
Acknowledgement
This project was supported by the Hong Kong Research Grants Council under Grants 11219324, 11219422, and 11202320.
Supplementary Material
Appendix 0.A Supplement of ablation study
In this section, we further discuss the effectiveness of some components of our method, as well as the robustness of the input data.
0.A.0.1 Keyframe selection and template surface learning.
Compared to directly setting the middle frame as a keyframe, Fig. 10 shows that the selected keyframe by our strategy is closer to the average shape. After learning the template surface, we also show the results of only using the fine stage (Fig. 10 (b)) and using the coarse-to-fine stage (Fig. 10 (c)). Based on Fig. 10 (c), we further show the enhanced template surface after the temporal reconstruction process.